OpenStudy (marcon):

Given a polynomial f(x), if (x + 3) is a factor, what else must be true? f(3) = 0 f(-3) = 0 f(0) = 3 f(0) = -3 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get so far on it?

OpenStudy (marcon):

I think I have to use the zero product property to figure it out

OpenStudy (diamondboy):

d factor of d function is -3 right?

jimthompson5910 (jim_thompson5910):

how would you use the zero product property?

OpenStudy (marcon):

Thats what I need help on @jim

OpenStudy (marcon):

Actually I'm not sure, we probably don't need to use the zero product property... I'm just confused on how to do this in general. Can you guys help me with the steps?

jimthompson5910 (jim_thompson5910):

ok let's say we had (x+3)(x-5) = 0 you would use the zero product property to get x+3 = 0 or x-5 = 0. Agreed?

OpenStudy (marcon):

mmhmm

OpenStudy (marcon):

What do we do next?

OpenStudy (marcon):

f(x) = (x + 3) = -3 f(-3) = 0 <-------- Correct? @jim_thompson5910

OpenStudy (marcon):

@Michele_Laino

jimthompson5910 (jim_thompson5910):

yes correct, if x+3 is a factor, then x = -3 is a root so f(-3) = 0

OpenStudy (marcon):

Thank you :D

jimthompson5910 (jim_thompson5910):

no problem